OpenStudy (anonymous):

Vector space V and subspace W. Let C(1) be the set of all continuously differentiable functions. V=C(1) and W consists of functions x(t) in C(1) satisfying the Integral w/ limits 4 to -2 of x(t)dt=0.

OpenStudy (turingtest):

\[\vec v=\int_4^{-2}x(t)dt\]this?

OpenStudy (anonymous):

\[\int\limits_{-2}^{4}x(t)dt=0\]

OpenStudy (turingtest):

\[\vec v_1=\int_{-2}^4x_1(t)dt\]\[\vec v_2=\int_{-2}^4 x_2(t)dt\]what is \[\vec v_1+\vec v_2\]?

OpenStudy (anonymous):

is it possible to even add those? I guess it has to be possible, but how do you do it since x(t) can be any function

OpenStudy (turingtest):

well you could imagine some examples\[x_1(t)=t^2-t\]\[x_2(t)=t^3\]

OpenStudy (turingtest):

what is\[\int_a^b t^2-tdt+\int_a^b t^3dt\]?

OpenStudy (zarkon):

@zonazoo what is your question?

OpenStudy (turingtest):

\[=\int_a^bt^3+t^2-tdt\]right?

OpenStudy (anonymous):

I have to show W is in the subspace of V

OpenStudy (turingtest):

you mean that you have to show that W is a subspace of V, not "in" the subpace

OpenStudy (turingtest):

subspace*

OpenStudy (anonymous):

yes that W is a subspace of V, sorry, trying to feed the kid and type at the same time.

OpenStudy (anonymous):

we can see examples, but i can't show it using examples can i

OpenStudy (turingtest):

\[\int_a^bf(t)dt+\int_a^bg(t)dt=\int_a^bf(t)+g(t)dt\]this should ring a bell

OpenStudy (zarkon):

clearly \(W\subseteq C(1)\) and \(W\) is nonempty (since \(\vec{0}\in W\)) now just show linearity

OpenStudy (anonymous):

how do I know that it = 0 though? because in order to be a subspace under addition it has to = 0 right?

OpenStudy (turingtest):

\[\vec u=\int_a^b x(t)dt=0\]\[\vec v=\int_a^b y(t)dt=0\]so\[\vec u+\vec v=?\]

OpenStudy (turingtest):

\[x(t),y(t)\in C_1\]

OpenStudy (anonymous):

let say for example instead of = 0 that it was = 1, if that it was the case then it wouldn't be a subspace right?

OpenStudy (turingtest):

right

OpenStudy (anonymous):

so really its more like anything except =0 it wouldn't be.

OpenStudy (turingtest):

in this case yes, but you should be able to show why.

OpenStudy (anonymous):

fantastic. not very comfortable with these as continuous functions. but it makes much more sense now. thank you. only a few more to go.

OpenStudy (turingtest):

you got it? welcome then

OpenStudy (anonymous):

yes I believe so.